Suppose a researcher is interested in the number of good versus bad dreams that students have during final exam week. The researcher states that p = 0.52 that a student will have a bad dream during final exam week. (a) What type of probability distribution is appropriate for these data? binomial distribution negatively skewed distribution bimodal distribution positively skewed distribution (b) If a professor has 50 students in his class, then how many students should he expect to have bad dreams during final exam week? ____students (c) Assuming complementary outcomes, what is the probability (q) that a student will have a good dream? q =_____
Answer #1

Answer

(A) BInomial distribution is correct answer because we have only two possible outcome (good or bad dreams) with fixed number of trails n and each trial is independent. So, Binomial distribution is correct answer

(B) Given that sample size n = 50

and probability p = 0.52

We know that Expected value E[x] = n*p

= 50*0.52

= 26

(C) Assuming complementary outcomes

then q = 1- p

where p = 0.52

so, probability of a good dream q = 1- 0.52

= 0.48